# A Simulated Annealing-Based Multiobjective Optimization Algorithm for Minimum Weight Minimum Connected Dominating Set Problem

### v1

@@ I Introduction
One should mention that all previous approaches have something in common: they optimize only a single objective function such as minimizing the size of the connected dominating set (CDS) and minimizing its total weight for MCDS and MWCDS, respectively. To the best of our knowledge, the first approach in the literature that considered these two objectives together was presented in [19]. The authors of this study firstly defined the minimum weight minimum connected dominating set problem (MWMCDS) of which the aim is to minimize simultaneously the size and the total weight of generated CDS. Then, they proposed a multiobjective genetic algorithm (MOGA) based on scalarization model to deal with MWMCDS problem.
In this work, we propose to solve the MWMCDS problem with a so-called multiobjective greedy simulated annealing (MGSA) algorithm. The concept of scalarization is used to provide a trade-off between size and total weight and a greedy heuristic is needed to seed simulated annealing with good initial solutions and generating neighbors.
The remainder of the paper is organized as follow: Section II describes the problem. Section III discusses the proposed MGSA algorithm. The experimental results are given in Section IV. Section V concludes the paper.
@@ II Problem statement
III Multiobjective Greedy Simulated Annealing algorithm MGSA for MWMCDS problem
In the following we describe the main components of MGSA algorithm that we developed with the aim of obtaining high quality solutions to the MWMCDS problem. As mentioned before, MGSA is a multiobjective greedy simulated annealing algorithm based on scalarization.
@@ III-D Multiobjective simulated annealing framework
Simulated Annealing (SA) [20] is a well known metaheuristic approach that has been applied successfully to a large number of combinatorial optimisation problems. SA is both a single-solution based algorithm and exploitation oriented. Our algorithm named MGSA is a standard SA algorithm improved by starting with a good initial solution based on the greedy heuristic previously defined (see Section III-B) and the neighbors are also generated either greedily or randomly. Moreover, MGSA uses a useful calculation method for changing temperature. These improvments ensure the computational efficiency and improve the quality of obtained solutions. A high level description of MGSA algorithm is illustrated in Alg. 1. The generate_initial_solution ($sol\_size$) procedure (see Alg. 2) generates $sol\_size$ solutions for the initial solution pool. Out of these solutions, one is generated greedily by applying generate_greedy() procedure (see Alg. 3) while the others are generated randomly using generate_random() procedure (see Alg. 4).
@@
Algorithm 1 MGSA for the MWMCDS problem
@@ IV EXPERIMENTAL EVALUATION
The proposed algorithm MGSA was implemented using C++ language. The experimental results were obtained on a PC with an Intel Core i5-1135G7 2.40GHz processor and 8.0 GB of memory.
Its performance was compared against two recent algorithms from the literature, namely MOGA [19] and mcds [21]. The results of the two approaches are reproduced from [19]. MGSA was evaluated on the same benchmark set introduced in [19] where each instance consists of a simple undirected edge-weighted graph modeling a data transfer system where every vertex transfers data at instant $t$ with probability $P_{t}$ and this data can be dropped with probability $P_{d}$. The distance traveled by the data in the transfer is represented indirectly by the weight, which expressed by the energy consumed during the travel. Thus, the energy consumed in case of successful transfer is equal to the distance traveled by the data, and in the case of failed transfer is equal to half of distance traveled. The energy consumed in transfer of all data in the network is represented by energy consumption.
The performance of MGSA against MOGA and mcds with respect to energy consumption for 100 instances of data transfer is shown in Fig. 2. It can be seen that MGSA consumes the least amount of energy in all cases compared with mcds and MOGA.
Fig. 3 represents the number of dominators vertices produced by MGSA, MOGA and mcds for different networks. Here too it can be seen that MGSA performs better than mcds in all cases by producing CDS of minimal size. Compared to MOGA, MGSA give the same results in 4 cases (n= 30, n=40, n=60 and n=70), in all other cases MGSA performs the best.
From previous results, it can be deduced that the proposed algorithm MGSA generates more useful solutions for the MWMCDS problem than mcds and MOGA.
Figure 2: Energy consumption in MGSA, MOGA and mcds
Figure 3: Size of CDS in MGSA, MOGA and mcds
@@ V Conclusion
In this work we have proposed a new approach called MGSA for tackling the minimum weight minimum connected dominating set problem in wireless sensor networks. Two objectives are involved together, namely the weight and the size of the connected dominating set. MGSA integrates a simulated annealing algorithm, which is seeded by a greedy constructive heuristic and employed in two hybridization models. The performance of the proposed algorithm was assessed and compared against a recent multiobjective genetic algorithm. For future work, it would be interesting to use Pareto optimization techniques instead of scalarization that reduces a multiobjective optimization problem to single objective optimization problem by a linear combination.

### v2

# Multiobjective Simulated Annealing Algorithm for Minimum Weight Minimum Connected Dominating Set Problem

@@ I Introduction
One should mention that all previous approaches have something in common: they optimize only a single objective function such as minimizing the size of the connected dominating set (CDS) and minimizing its total weight for MCDS and MWCDS, respectively. To the best of our knowledge, the only existing approach in the literature that considered these two objectives together was presented in [19]. The authors of this study firstly defined the minimum weight minimum connected dominating set problem (MWMCDS) of which the aim is to minimize simultaneously the size and the total weight of generated CDS. Then, they proposed a multiobjective genetic algorithm (MOGA) based on scalarization model to deal with MWMCDS problem.
In this work, we propose to solve the MWMCDS problem with a so-called multiobjective greedy simulated annealing (GSA) algorithm. The concept of scalarization is used to provide a trade-off between size and total weight and a greedy heuristic is needed to seed simulated annealing with good initial solutions and generating neighbors.
The remainder of the paper is organized as follow: Section II describes the problem. Section III discusses the proposed GSA algorithm. The experimental results are given in Section IV. Section V concludes the paper.
@@ II Problem statement
III Multiobjective Greedy Simulated Annealing algorithm GSA for MWMCDS problem
In the following we describe the main components of GSA algorithm that we developed with the aim of obtaining high quality solutions to the MWMCDS problem. As mentioned before, GSA is a multiobjective greedy simulated annealing algorithm based on scalarization.
@@ III-D Multiobjective simulated annealing framework
Simulated Annealing (SA) [20] is a well known metaheuristic approach that has been applied successfully to a large number of combinatorial optimisation problems. SA is both a single-solution based algorithm and exploitation oriented. Our algorithm named GSA is a standard SA algorithm improved by starting with a good initial solution based on the greedy heuristic previously defined (see Section III-B) and the neighbors are also generated either greedily or randomly. Moreover, GSA uses a useful calculation method for changing temperature. These improvments ensure the computational efficiency and improve the quality of obtained solutions. A high level description of GSA algorithm is illustrated in Alg. 1. The generate_initial_solution ($sol\_size$) procedure (see Alg. 2) generates $sol\_size$ solutions for the initial solution pool. Out of these solutions, one is generated greedily by applying generate_greedy() procedure (see Alg. 3) while the others are generated randomly using generate_random() procedure (see Alg. 4).
@@
Algorithm 1 GSA for the MWMCDS problem
@@ IV EXPERIMENTAL EVALUATION
The proposed algorithm GSA was implemented using C++ language. The experimental results were obtained on a PC with an Intel Core i5-1135G7 2.40GHz processor and 8.0 GB of memory.
Its performance was compared against two recent algorithms from the literature, namely MOGA [19] and mcds [21]. The results of the two approaches are reproduced from [19]. GSA was evaluated on the same benchmark set introduced in [19] where each instance consists of a simple undirected edge-weighted graph modeling a data transfer system where every vertex transfers data at instant $t$ with probability $P_{t}$ and this data can be dropped with probability $P_{d}$. The distance traveled by the data in the transfer is represented indirectly by the weight, which expressed by the energy consumed during the travel. Thus, the energy consumed in case of successful transfer is equal to the distance traveled by the data, and in the case of failed transfer is equal to half of distance traveled. The energy consumed in transfer of all data in the network is represented by energy consumption.
The performance of GSA against MOGA and mcds with respect to energy consumption for 100 instances of data transfer is shown in Fig. 2. It can be seen that GSA consumes the least amount of energy in all cases compared with mcds. Compared to MOGA, the energy requirements of GSA are lesser in all cases except the scenario ($n=100$) where both algorithms give similar results.
Fig. 3 represents the number of dominators vertices produced by GSA, MOGA and mcds for different networks. Here too it can be seen that GSA performs better than mcds in all cases by producing CDS of minimal size. Compared to MOGA, GSA give the same results in 4 cases (n= 30, n=40, n=60 and n=70), in all other cases GSA performs the best.
From previous results, it can be deduced that the proposed algorithm GSA generates more useful solutions for the MWMCDS problem than mcds and MOGA.
Figure 2: Energy consumption in GSA, MOGA and mcds
Figure 3: Size of CDS in GSA, MOGA and mcds
@@ V Conclusion
In this work we have proposed a new approach called GSA for tackling the minimum weight minimum connected dominating set problem in wireless sensor networks. Two objectives are involved together, namely the weight and the size of the connected dominating set. GSA integrates a simulated annealing algorithm, which is seeded by a greedy constructive heuristic and employed in two hybridization models. The performance of the proposed algorithm was assessed and compared against a recent multiobjective genetic algorithm. For future work, it would be interesting to use Pareto optimization techniques instead of scalarization that reduces a multiobjective optimization problem to single objective optimization problem by a linear combination.